### v1

# FedRef: Communication-Efficient Bayesian Fine Tuning with Reference Model

Abstract
Federated learning(FL) is used for distributed scenarios to train artificial intelligence(AI) models while ensuring users’ privacy. In federated learning scenario, the server generally never knows about users’ data. This type of concept makes the AI training process efficient in terms of data privacy. However, regarding model performance, federated AI models may not sufficiently satisfy AI users’ expectations. Furthermore, AI users have a wide range of different needs. It is not easy to satisfy the whole users needs. These types of issues can be addressed through AI model optimization, fine-tuning, or personalization to achieve optimal model performance. To address model optimization challenges, we propose reference model-based federated learning for optimal fine-tuning, which overcomes catastrophic forgetting in each round. This method is derived from Bayesian parameter-efficient transfer learning, which includes an optimal proximal term and enables overcoming the catastrophic forgetting issue in each round by utilizing a reference model that incorporates previous model parameters. As a result, this method achieves both high model performance and low computing cost.
Index Terms:
Privacy Preserving Machine Learning, Federated Learning, Bayesian Optimization, Transfer Learning, Fine-Tune.
I Introduction
Recently, in the AI industry, federated learning (FL) has been proposed as a prominent solution to protect users’ data privacy while allowing collaborative model training between independent affiliations. Users’ data are fundamentally protected because clients share only updated local model parameters (resulting from local training using their local data) and other non-private values with the server or other clients. The server then computes the global model by aggregating the local models received by the clients. However, as highlighted in [1], FL solutions face numerous challenges. Our focus is on optimizing model performance and decreasing computing costs on client devices.
  1. 1.
    Model performance optimization in FL scenarios.
  2. 2.
    Decreasing computing cost and energy consumption.
  3. 3.
    How to protect the model from malicious users.
In the related work section, we introduce existing approaches (FedAvg, FedProx, FedOpt, Bayesian Transfer Learning) for model optimization in FL scenarios, along with their limitations regarding client computing cost and catastrophic forgetting [2][3] [4] in each round. We then introduce our FedRef: a communication-efficient Bayesian fine-tuning approach with a reference model. This concept overcomes catastrophic forgetting in each round by integrating previous model features into the maximum a posteriori (MAP) problem.
In existing work, the MAP problem was applied as an objective function for optimizing model transfer learning between a target model with pre-trained parameters and a source model. On the other hand, traditional model optimization in FL scenarios used a proximal term to minimize deviation from the global model. However, the proximal term can be denoted as a MAP problem. In this paper, we define the MAP problem as an objective function that is updated with aggregated models from previous rounds only on the server-side, thereby minimizing clients’ computing cost.
In our experiments, we applied FedRef on MNIST, CIFAR10, FeTs2022 data partitioning from distributed condition for non-IID(Independent and Identically Distributed) setting. We defined each baseline and then analyzed model performances, as baseline models differ for each dataset. Especially, on the FeTs2022, the data is derived in medical imaging area for brain tumor segmentation. In this case, the client-side can be a hospital, as shown in Fig. 1.
Figure 1: Basic federated learning system.
As a result, our FedRef achieves high model performance, despite minimizing clients’ computing cost.
II Related Work
II-1 FedAvg
Communication-efficient learning of deep networks from decentralized data [1], they proposed typical system of federated learning concept. The paper, which introduced FedAvg, demonstrated higher model performance than FedSGD on the MNIST classification task under both IID and non-IID conditions using the Federated Averaging method. Federated Averaging is based on the stochastic gradient descent (SGD) concept [5], which aggregates client models weighted by the number of samples from each client, as shown in equation (1).
$\theta_{r+1}=\sum_{i=1}^{K}\frac{n_{i}}{n}\theta_{i,r}$ (1)
$r$: round of repeating server routine, $\theta_{i}$ : $i$ index client’s model parameter, $n_{i}$: $i$ index client’s number of samples, $n$: total number of clients’ samples. We also selected this type of aggregation, but it is easy to apply other weighted aggregation concepts to our FedRef solution. On the client side, local training is represented by equation (2).
$\theta_{i}^{(s+1)}=\theta_{i}^{(s)}-\eta\nabla F_{i}(\theta_{i}^{(s)};\mathcal{B}_{i}^{(s)})$ (2)
$s$ : number of local steps for local training, $\eta$: learning rate, $\mathcal{B}_{i}^{(s)}$: mini-batch sampled from $i$ index client local dataset on step $s$, $F_{i}(\theta_{i}^{(s)};\mathcal{B}_{i}^{(s)})$: local loss function evaluated on a mini-batch. The local training in our FedRef client update is based on the SGD solution. In algorithm 1, it show whole FedAvg process flow.
Algorithm 1 FedAvg: Communication-efficient learning of deep networks from decentralized data
1:Initialize global model weights $\theta_{0}$
2:for each round $r=1,2,\dots,R$ do
3:     Server selects of clients $\mathcal{S}_{r}\subseteq\{1,\dots,K\}$
4:     for all client $k\in\mathcal{S}_{r}$ in parallel do
5:         Client $k$ receives $\theta_{r}$ from the server
6:         Client $k$ performs local training:
7:          $\theta_{r+1}^{k}\leftarrow\text{ClientUpdate}(k,\theta_{r})$
8:     end for
9:     Server aggregates the client updates:
10:      $\theta_{r+1}\leftarrow\sum_{k\in\mathcal{S}_{r}}\frac{n_{k}}{n_{\mathcal{S}_{r}}}\theta_{r+1}^{k}$
11:      $n_{\mathcal{S}_{r}}=\sum_{k\in\mathcal{S}_{r}}n_{k}$
12:end for
II-2 FedProx
Federated Optimization in Heterogeneous Networks [6], they proposed equation (3) named ”Proximal Term” to objective function for ensuring that local models do not deviate significantly from the global model. This concept enables optimizing the model to account for data skew and heterogeneity. However, it does not consider clients’ computing costs, as clients must compute the proximal term and send it to the server.
$\min_{\theta_{k}}\left(\mathbf{E}_{\mathcal{D}_{k}}\left[\mathcal{L}(\theta_{k},\mathcal{D}_{k})\right]+\frac{\mu}{2}\|\theta_{k}-\theta_{\text{global}}\|^{2}\right)$ (3)
$k$ : a client index, $\theta_{k}$ : a client $k$ local model parameter, $\mathcal{L}(\theta_{k},\mathcal{D}_{k})$ : client $k$ local data $\mathcal{D}_{k}$ cost function, $\theta_{\text{global}}$ : global model parameter, $\mu$: hyperparameter for proximal term strength, $\|\theta_{k}-\theta_{\text{global}}\|^{2}$: L2(Euclidean) distance between local model parameter and global model parameter. The complete FedProx Algorithm 2 is shown below.
Algorithm 2 FedProx: Federated Optimization in Heterogeneous Networks
1:Initialize global model weights $\theta_{0}$
2:for each round $r=1,2,\dots,R$ do
3:     Server selects clients $\mathcal{S}_{r}\subseteq\{1,\dots,K\}$
4:     for all client $k\in\mathcal{S}_{r}$ in parallel do
5:         Client $k$ receives $\theta_{r}$ from the server
6:         Client $k$ performs local training:
7:          $\theta_{r+1}^{k}\leftarrow\text{ClientUpdate\_Prox}(k,\theta_{r},\mu)$
8:     end for
9:     Server aggregates the client updates:
10:      $\theta_{r+1}\leftarrow\sum_{k\in\mathcal{S}_{r}}\frac{n_{k}}{n_{\mathcal{S}_{r}}}\theta_{r+1}^{k}$
11:      where $n_{k}$ is the number of samples at client $k$, and $n_{\mathcal{S}_{r}}=\sum_{k\in\mathcal{S}_{r}}n_{k}$
12:end for
In the server side, they just aggregate and return model parameter to clients. This led to the issue of concentrated computing costs on the client. Therefore, on the client side, Algorithm 3 introduces local training with a proximal term, which increases the client computing cost.
Algorithm 3 ClientUpdate_Prox($k$, $\theta_{r}$, $\mu$)
1:Initialize local model $\theta\leftarrow\theta_{r}$
2:for each local epoch $e=1,2,\dots,E$ do
3:     for each minibatch $b$ in local data $\mathcal{D}_{k}$ do
4:         Compute gradient:
5:          $g\leftarrow\nabla f_{k}(\theta;b)+\mu(\theta-\theta_{r})$
6:         Update local model:
7:          $\theta\leftarrow\theta-\eta g$
8:     end for
9:end for
10:return $\theta$
II-3 FedOpt
Adaptive Federated Optimization [7], they introduced and included optimization concept merged with (FedADAGRAD, FedADAM, FedYOGI). They included each different specific proximal terms. Following FedADAGRAD equation (4):
$g_{r}=\frac{1}{K}\sum_{k=1}^{K}\Delta\theta_{r}^{(k)},$
$G_{r}=G_{r-1}+g_{r}^{2},$
$\theta_{r+1}=\theta_{r}-\eta\cdot\frac{g_{r}}{\sqrt{G_{r}}+\tau}$ (4)
$g_{r}$: averaged update from all participating clients, $G_{r}$: accumulated sum of squared gradients (maintained on the server), $\tau$: small constant for numerical stability (e.g., $10^{-6}$). They introduced optimization of local models to a variant that considers global data distribution with two types of different parameters. Following FedADAM equation (II-3):
$m_{r}=\beta_{1}m_{r-1}+(1-\beta_{1})g_{r},$
$v_{r}=\beta_{2}v_{r-1}+(1-\beta_{2})g_{r}^{2},$
$\hat{m}_{r}=\frac{m_{r}}{1-\beta_{1}^{r}},\,\hat{v}_{r}=\frac{v_{r}}{1-\beta_{2}^{r}},$
$\theta_{r+1}=\theta_{r}-\eta\cdot\frac{\hat{m}_{r}}{\sqrt{\hat{v}_{r}}+\tau}$ (5)
$m_{r}$: first moment vector (mean of gradients), $v_{r}$: second moment vector (uncentered variance of gradients), $\hat{m}_{r},\hat{v}_{r}$: bias-corrected moment estimates, $\beta_{1},\beta_{2}$: decay rates for first and second moments (e.g., $\beta_{1}=0.9,\beta_{2}=0.999$). The key term in FedADAM also denotes the variance between aggregated gradients and uncentered model parameters. Finally following FedYOGI equation (II-3):
$v_{t}=v_{r-1}-(1-\beta_{2})\cdot sign(v_{r-1}-g_{r}^{2})\cdot g_{r}^{2},$
$m_{r}=\beta_{1}m_{r-1}+(1-\beta_{1})g_{r},$
$\hat{m}_{r}=\frac{m_{r}}{1-\beta_{1}^{r}}\,\hat{v}_{r}=\frac{v_{r}}{1-\beta_{2}^{r}},$
$w_{r+1}=w_{r}-\eta\cdot\frac{\hat{m}_{r}}{\sqrt{\hat{v}_{r}}+\tau}$ (6)
sign()($\cdot$)( ⋅ ): Element-wise sign function. This paper introduced various proximal term concepts (FedADAGRAD, FedADAM, FedYOGI) that are automatically selected based on specific conditions. For details, we present the complete FedOpt in Algorithm4:
Algorithm 4 FedOpt: Adaptive Federated Optimization
1:Input: $\theta_{0}$, CLIENTOPT, SERVEROPT
2:
3:for $r=0,\dots,R-1$ do
4:     Sample a subset $\mathcal{S}$ of clients
5:     $\theta_{r,0}^{i}=\theta_{r}$
6:     for each client $i\in\mathcal{S}$ in parallel do
7:         for $k=0,\dots,K-1$ do
8:              Compute $g_{r,k}^{i}$ of $\nabla F_{i}(\theta_{r,k}^{i})$
9:              $\theta_{r,k+1}^{i}=\texttt{CLIENTOPT}(\theta_{r,k}^{i},g_{r,k}^{i},\eta_{l},r)$
10:         end for
11:         $\Delta_{r}^{i}=\theta_{r,K}^{i}-\theta_{r}$
12:     end for
13:     $\Delta_{r}=\frac{1}{|\mathcal{S}|}\sum_{i\in\mathcal{S}}\Delta_{r}^{i}$
14:     $\theta_{r+1}=\texttt{SERVEROPT}(\theta_{r},-\Delta_{r},\eta,r)$
15:end for
These methods also required client-side computing for the proximal term, which needed to be optimized. Regarding state-of-the-art strategies, recent similar research (FedDyn [8]) has been proposed, which adaptively selects the optimal strategy. However, both FedOpt and FedDyn needed client side computing costs for specific proximal terms.
II-4 Bayesian Transfer Learning
Bayesian parameter-efficient fine-tuning for overcoming catastrophic forgetting [9], This paper proposed a Bayesian transfer learning concept for overcoming the catastrophic forgetting issue and formulated the optimal model fine-tuning cost as a Maximum A Posteriori (MAP) problem in equation(7).
$\theta^{*}=\arg\max_{\theta}\left[\log p(D_{B}|\theta)+\log p(\theta|D_{A})\right]$ (7)
$D_{A}$: data used for pretraining, $D_{B}$: data used for downstream fine-tuning. In equation (7), $\log p(\theta|D_{A})$ can be approximated by Laplace Approximation [10] to yield equation (8).
$\log p(\theta|D_{A})\approx f(\theta_{0})+\tfrac{1}{2}(\theta-\theta_{0})^{\top}H(\theta-\theta_{0})$ (8)
$H$: negative Hessian of the log-posterior, approximated as the FIM(Fisher Information Matrix). Finally, they defined the objective function for the optimization of transfer learning in equation (9).
$\mathcal{L}(\theta)=\underbrace{L_{B}(\theta)}_{\text{task loss}}+\lambda\,(\theta-\theta_{0})^{\top}F\,(\theta-\theta_{0})$ (9)
$\lambda$: regularization strength, $F$: approximation of the Fisher Information Matrix or Hessian. According to this paper, to decrease computing cost and simplify the FIM matrix to an identity matrix, despite decreasing the capability of detecting model attention (which has key points affecting model performance), equation (9) can be simplified to equation (10).
$\mathcal{L}_{\text{l2-sp}}(\theta)=L_{B}(\theta)+\lambda\sum_{k}(\theta_{k}-\theta_{0,k})^{2}$ (10)
Despite being simplified, equation (10) has been shown to increase the rate of preventing catastrophic forgetting in transfer learning [11][12]. For other Hessian metric assumptions, they described Kronecker-factored approximation[13][14], which also considers the covariance between parameters within each layer. As a result, this paper introduced optimization of transfer learning between only two models: one a pre-trained model intended for transfer, and the other a target model. However, in an FL scenario, the server must aggregate and optimize many different client models. Therefore, we need a definition of a new MAP cost problem for FL scenarios.
III FedRef: Communication-Efficient Bayesian Fine Tuning with Reference Model
For optimal model performance and low client computing cost, we proposed FedRef: communication-efficient Bayesian fine tuning with reference model which overcome catastrophic forgetting by inferring of previous rounds model. We carefully defined the model proximal term as a MAP problem for FL scenarios in equation (11) derived from the Bayesian fine-tuning paper[9].
$\theta^{*}=\arg\max_{\theta}[\log p(\theta|D_{ref})+\log p(D_{1}|\theta)+$
$logp(D_{2}|\theta)+\dots+\log p(D_{K}|\theta)]$ (11)
Selected client numbers $k\in[1,2,3,..,K]$, where $K$ is the total number of selected clients. $D_{ref}$: represents synthetic data of a reference model that incorporates features from previous rounds. $D_{ref}$ is defined solely for explaining our optimal MAP problem (III). This concept enables overcoming catastrophic forgetting in each round by integrating previous round features into the MAP problem and then optimizing the integrated MAP value. Finally, the objective function derived in (11) can be expressed by combining equations (8, 9, 10) to form equation (III).
$L_{ref}(\theta^{1},\theta^{2},\forall F)=\sum_{k}^{K}F_{k}\cdot\mathrm{diag}(W_{1},\dots,W_{K})$
$+\lambda\sum_{i}(\theta^{1}_{i}-\theta^{1}_{0,i})^{2}+\lambda\sum_{i}(\theta^{2}_{i}-\theta^{2}_{0,i})^{2}$ (12)
In equation (III), the constant value $\sum_{k}^{K}F_{k}$ denotes the sum of client losses. The diagonal matrix $\mathrm{diag}(W_{1},\dots,W_{K})$ represents aggregation weights (e.g. $\frac{n_{i}}{n}$). $\sum_{i}(\theta_{i}-\theta_{0,i})^{2}$ signifies $L_{2}$ regularization of model $\theta-\theta_{0}$. Regarding parameter requirements, only clients’ losses are needed for our FedRef concept. Only on the server side, model optimization is performed, which can decrease both client computing cost and communication cost. In the FedRef, $\theta^{2}$ should be set as the reference model $\theta_{ref}$. For detail on $\theta_{ref}$, we defined the reference model as shown in equation 13.
$\theta_{r+1}^{2}=\text{A}(\theta^{1}_{r-p},\theta^{1}_{r-p+1},\dots,\theta^{1}_{r})$ (13)
In the equation (13), $p$: the number of selected subset of previous aggregation model which can be set heuristically and A function should be as following a equation 14.
$\text{A}(\forall\theta)=\sum_{i=1}^{p}\frac{1}{p}\theta^{1}_{r+1-i}$ (14)
Experimentally, a suitable value for $p$ was analyzed to be 5. For the whole algorithm, we denoted FedRef at following algorithm5:
Algorithm 5 FedRef: Communication-Efficient Bayesian Fine Tuning with Reference Model
1:Client:
2:for epoch do
3:     for batch do
4:   $\theta_{i}^{(s+1)}=\theta_{i}^{(s)}-\eta\nabla F_{i}(\theta_{i}^{(s)};\mathcal{B}_{i}^{(s)})$
5:     end for
6:end for
7:send $\theta$ , $F(\theta^{(s)};\mathcal{B}^{(s)})$ to Server
8:
9:Server:
10:for round do
11:$K$-client select
12:$A_{r}$leftarrow $\text{Aggregation}(\theta_{1},\theta_{2},\theta_{3}...\theta_{K})$
13:$R_{r}$leftarrow $\text{Aggregation}(A_{r-p}...,A_{r-1},A_{r})$
14:$\theta$leftarrow Bayesian $\text{Bayesian Optimization}(A_{r},R_{r},\forall F(\theta^{(s)};\mathcal{B}^{(s)}))$
15: broadcast $\theta$ to Clients
16:end for
17:
[…]
19:$\theta_{r+1}=\sum_{i=1}^{K}\frac{n_{i}}{n}\theta_{i,r+1}$
20:return $\theta$
21:
22:Bayesian Optimization($\theta^{1},\theta^{2}$):
23:$\theta^{1}\leftarrow\theta^{1}-\eta\times\nabla{L}_{ref}(\theta^{1},\theta^{2})$
24:return $\theta^{1}$
25:
[…]
IV Experiment
IV-A Experiment Setting
IV-A1 Global
We developed a federated learning system using the Flower framework [15]. For detailed experimental conditions, table I presents the global computing environment settings.
TABLE I: Experiment Environment
Setting Specification
FL Framework Flower: a friendly federated learning framework
Language Python
Operation System Linux 24.04 LTS
GPU Nvidia RTX 4090
Tools Visual studio code
Within the Python language environment, the Flower framework enables easy construction of federated learning servers and clients from high-level code, based on Object-Oriented Programming (OOP) and Functional Programming (FP), for privacy-preserving machine learning strategies(e.g. FedAvg, FedProx, FedOpt, etc.).
IV-A2 MNIST
The MNIST (Modified National Institute of Standards and Technology) dataset[16] is one of the most widely used datasets in the field of machine learning and computer vision. It serves as a benchmark for evaluating the performance of various algorithms, particularly in the area of handwritten digit recognition. We discuss federated leaning models that have been applied to this dataset. Each client has a partitioning rate of [50 to 90%] for train samples, and this rate is randomly varied for the non-IID setting. For the model baseline, MobileNet [17] is applied for our MNIST classification task experiment.
IV-A3 CIFAR10
The CIFAR-10 (Canadian Institute for Advanced Research) dataset[18] is one of the most prominent and widely used image datasets in machine learning and computer vision research. It consists of 60,000 32x32 color images across 10 distinct classes, making it a popular benchmark for evaluating the performance of classification algorithms, particularly in the realm of image recognition. The data partitioning mechanism for this dataset is the same as the MNIST setting. For the model baseline, ResNet [19] is applied for CIFAR10 classification task.
IV-A4 FeTs2022
The FeTs2022 dataset refers to the dataset used in the Federated Tumor Segmentation (FeTS) Challenge 2022, which focused on brain tumor segmentation using federated learning approaches. The data includes multimodal brain MRI scans of glioma patients. Each patient folder typically has the four MRI modalities (T1-weighted, T1-weighted post-contrast, T2-weighted, T2 Fluid Attenuated Inversion Recovery). The data labels include segmentations of three tumor sub-regions (Enhancing tumor, Tumor core, Whole tumor). The data can be downloaded at official Synapse site [20]. We perform the segmentation task for evaluating FedRef on FeTs2022 data which have specific non-IID data partitioning at Fig. 2.
Figure 2: Data partitioning of the FeTS2022 dataset.
IV-B Result
Figure 3: Graph of centralized loss on each strategy and data
TABLE II: Performance comparison on segmentation task from FeTs2022
Data Strategy mIoU Dice-score Hausdorff 95
Fed-Avg
[…]
FedOpt
FedRef
V Challenge
FedRef still faces many challenges, including the selection of hyperparameters (client-server side learning rate, a constant value representing the total number of selected previous round models) and protection against malicious users and model inference attacks. Regarding assumptions about the FIM matrix, there are various equations. For future work, we should analyze various FIM matrices and define a more proper proximal term.
VI Conclusion
For optimizing model training in FL scenarios, existing works have proposed various proximal terms that act as regularization terms in the objective function, preventing the local model from drifting too far from the global model. However, they did not solve the catastrophic forgetting issue in each round, and the proximal term had to be computed on the client side, which increased the client’s computing cost. Our FedRef, a communication-efficient Bayesian fine-tuning approach with a reference model, solves these issues by applying server-side only optimization, including a proximal term that infers previous global features. This concept leads to low client-side computing cost and high model optimization performance. As a result, FedRef contributes to model training optimization in federated learning conditions.
Acknowledgment
The preferred spelling of the word “acknowledgment” in America is without an “e” after the “g”. Avoid the stilted expression “one of us (R. B. G.) thanks $\ldots$”. Instead, try “R. B. G. $\ldots$”. Put sponsor acknowledgments in the unnumbered footnote on the first page.

### v2

@@
 To address model optimization challenges, we propose reference model-based federated learning for optimal fine-tuning, which overcomes catastrophic forgetting in each round.
-This method is derived from Bayesian parameter-efficient transfer learning, which includes an optimal proximal term and enables overcoming the catastrophic forgetting issue in each round by utilizing a reference model that incorporates previous model parameters.
-As a result, this method achieves both high model performance and low computing cost.
+This method is derived from Bayesian parameter-efficient transfer learning, which includes an optimal proximal term and utilizes a reference model that incorporates previous model parameters.
+As a result, this method achieves both high model performance and clients’ low computing cost.
 Index Terms:
@@
     How to protect the model from malicious users.
-In the related work section, we introduce existing approaches (FedAvg, FedProx, FedOpt, Bayesian Transfer Learning) for model optimization in FL scenarios, along with their limitations regarding client computing cost and catastrophic forgetting [2][3] [4] in each round.
+In the related work section, we introduce existing approaches (FedAvg, FedProx, FedOpt, Bayesian Transfer Learning) for model optimization in FL scenarios, along with their limitations regarding client computing cost and catastrophic forgetting [2][3] [4] issue.
 We then introduce our FedRef: a communication-efficient Bayesian fine-tuning approach with a reference model.
@@
 In this paper, we define the MAP problem as an objective function that is updated with aggregated models from previous rounds only on the server-side, thereby minimizing clients’ computing cost.
-In our experiments, we applied FedRef on MNIST, CIFAR10, FeTs2022 data partitioning from distributed condition for non-IID(Independent and Identically Distributed) setting.
+In our experiments, we applied FedRef on MNIST, DRIVE, FeTs2022, BRATS data partitioning from distributed condition for non-IID(Independent and Identically Distributed) setting.
 We defined each baseline and then analyzed model performances, as baseline models differ for each dataset.
-Especially, on the FeTs2022, the data is derived in medical imaging area for brain tumor segmentation.
+Especially, on the FeTs2022 and BRATS, the dataset are derived in medical imaging area for brain tumor segmentation.
 In this case, the client-side can be a hospital, as shown in Fig.
 1.
-Figure 1: Basic federated learning system.
-As a result, our FedRef achieves high model performance, despite minimizing clients’ computing cost.
+As a result, our FedRef achieves high model performance, despite minimizing clients’ computing cost in medical area and general computer vision tasks.
+In the Result section, we represented models performance of each round by graph and specifically, the FeTs2022 task table shows specific evaluation metrics (mean Intersection of Union, Dice Coefficient, Hausdorff Distance) [5][6][7][8] which generally be used on medical segmentation tasks.
+Figure 1: Basic medical federated learning system.
 II Related Work
@@
 $s$ : number of local steps for local training, $\eta$: learning rate, $\mathcal{B}_{i}^{(s)}$: mini-batch sampled from $i$ index client local dataset on step $s$, $F_{i}(\theta_{i}^{(s)};\mathcal{B}_{i}^{(s)})$: local loss function evaluated on a mini-batch.
-The local training in our FedRef client update is based on the SGD solution.
 In algorithm 1, it show whole FedAvg process flow.
@@
 Finally following FedYOGI equation (II-3):
-$v_{t}=v_{r-1}-(1-\beta_{2})\cdot sign(v_{r-1}-g_{r}^{2})\cdot g_{r}^{2},$
+$v_{r}=v_{r-1}-(1-\beta_{2})\cdot sign(v_{r-1}-g_{r}^{2})\cdot g_{r}^{2},$
 $m_{r}=\beta_{1}m_{r-1}+(1-\beta_{1})g_{r},$
@@
 Regarding parameter requirements, only clients’ losses are needed for our FedRef concept.
-Only on the server side, model optimization is performed, which can decrease both client computing cost and communication cost.
+Only on the server side, model optimization is performed, which can decrease client computing cost.
 In the FedRef, $\theta^{2}$ should be set as the reference model $\theta_{ref}$.
@@
 $\text{A}(\forall\theta)=\sum_{i=1}^{p}\frac{1}{p}\theta^{1}_{r+1-i}$ (14)
-Experimentally, a suitable value for $p$ was analyzed to be 5. For the whole algorithm, we denoted FedRef at following algorithm5:
+Experimentally, a suitable value for $p$ was analyzed heuristically to be 3 to 5, but it should be considered about memory resources for saving parameters.
+For the whole algorithm, we denoted FedRef at following algorithm5:
 Algorithm 5 FedRef: Communication-Efficient Bayesian Fine Tuning with Reference Model
@@
 FedAvg, FedProx, FedOpt, etc.).
+The source code for our method is publicly available at [https://github.com/TaehwanY98/Fed-Ref]
 IV-A2 MNIST
@@
 For the model baseline, MobileNet [21] is applied for our MNIST classification task experiment.
-IV-A3 CIFAR10
-The CIFAR-10 (Canadian Institute for Advanced Research) dataset[18] is one of the most prominent and widely used image datasets in machine learning and computer vision research.
-It consists of 60,000 32x32 color images across 10 distinct classes, making it a popular benchmark for evaluating the performance of classification algorithms, particularly in the realm of image recognition.
-The data partitioning mechanism for this dataset is the same as the MNIST setting.
-For the model baseline, ResNet [19] is applied for CIFAR10 classification task.
-IV-A4 FeTs2022
+IV-A3 FeTs2022
 The FeTs2022 dataset refers to the dataset used in the Federated Tumor Segmentation (FeTS) Challenge 2022, which focused on brain tumor segmentation using federated learning approaches.
@@
 2.
+Baseline model is used to 3D U-Net.
+IV-A4 Drive
+The DRIVE dataset is a widely recognized resource in the field of retinal image analysis, particularly for blood vessel segmentation.
+It helps researchers study, diagnose, and manage various cardiovascular and ophthalmic diseases like diabetes and hypertension by analyzing morphological properties of retinal vessels such as length, width, tortuosity, and branching patterns.
+The dataset comprises 40 color fundus photographs.
+Seven of these images show signs of mild early diabetic retinopathy, representing abnormal pathological cases.
+All images were collected from a diabetic retinopathy screening program in the Netherlands [23].
 Figure 2: Data partitioning of the FeTS2022 dataset.
 IV-B Result
-Figure 3: Graph of centralized loss on each strategy and data
-TABLE II: Performance comparison on segmentation task from FeTs2022
+Figure 3: Graph of server-side centralized loss on each strategy and data.
+Figure 4: Graph of server-side centralized Dice-coefficient(DCE) and Hausdorff distance. Left: DCE, Right: HF95
+TABLE II: Performance comparison from Medical data
 Data Strategy mIoU Dice-score Hausdorff 95
-Fed-Avg
+Fed-Avg 16.78% 21.49% 15.29
 […]
-FedOpt
-FedRef
+FedOpt 16.67% 21.39% 12.54
+FedRef 16.81% 21.53% 12.13
+Fed-Avg 30.23% 45.95% 14.82
+[…]
+FedOpt 30.23% 45.95% 14.82
+FedRef 32.27% 48.17% 14.79
 V Challenge